### v1

@@ A Relationship to Biosphere B Spatial Distribution C Development D Highest Order E Smallest Scale
As we did in §2.1, we conduct a cross-consistency assessment for the value pairs in the technology factors matrix to eliminate any pairs that would be inconsistent. The use of the LLM is even more helpful in providing explicit concise arguments for and against cross-consistency for this second matrix, as there are 113 unique possible combinations of value pairs in Table 3; however, we still need to manually inspect and evaluate all the initial results and rationales and make corrections. The resulting technology cross-consistency matrix eliminates 50 of the 113 possible value pairs as inconsistent or unlikely, as shown in Figure 2.
Figure 2: Technology factors cross-consistency assessment. The 50 cells marked ‘x’ indicate inconsistent value pairs. The remaining 63 open cells indicate consistent value pairs.
With the cross-consistency assessment complete, the remaining value pairs in the technology factors cross-consistency matrix (Fig. 2) result in eleven viable scenarios, which are summarized in Table 4. To better imagine and comprehend these value combinations, we again use the Claude LLM to generate illustrative descriptions of these eleven technology factors scenarios for a thousand years into the future. The LLM scenario descriptions suggested that some scenarios show significant similarities to the extent that they could be grouped together. We choose to manually cluster the technology factors scenarios in an effort to further reduce the total number required for analysis.
@@ 2.3 Final Scenarios
We perform a compatibility assessment of the 33 viable global and technology scenarios to identify any combinations of global and technology factors scenarios that are inconsistent on the basis of their corresponding myths/metaphors that reflect their deeper civilizational foundations. We take an interactional theory stance on the immaterial and material manifestations of the future anthroposphere, captured by the global and technology factors: we assume that they must be mutually shaped and compatible from a civilizational perspective at their deepest cultural level, that is, at the level of collective metaphors and myths. We invoke the Claude LLM to initially illustrate this compatibility assessment, which we then manually review and correct as needed. The results of the compatibility assessment are shown in Figure. 3, which leaves 16 remaining combinations between the global and technology factors that are compatible at the myth/metaphor level of their underpinning civilizations.
Figure 3: Compatibility assessment between global factors scenarios and technology factors scenarios. The 17 cells marked ‘x’ indicate inconsistent pairings. The remaining 16 open cells indicate consistent parings.
@@ 5.1 Kardashev Scale Revisited
Technology exists to solve human biological and social problems, and technospheres evolve as the aggregate of these solutions. The methodology in this paper provides a way to link insights from the social sciences and humanities for understanding the complex dynamics of human societies into projections of the future that can inform the search for technosignatures. The underlying premise of this work is that human technology emerges from human needs and becomes intertwined with humans, which means that we cannot make meaningful projections about future technology without also thinking about how future developments would co-evolve with our human needs. This is one of the key insights missing from the projections by Kardashev (1964).
@@
We use the values in Table 9 to show approximate trajectories for energy use across the 1000-year timeline for all our scenarios in Figure 10. We draw a line or curve from the present-day energy use to the final value at the end of the timeline for each scenario. For some scenarios the value of energy use is lower than today (S4, S7, S8), which results from a prior period of growth before collapsing into the current state. Scenario S8 has a long-period oscillation with rapid growth followed by a collapse, currently midway through its third phase of growth. The three stable scenarios (S3, S5, S10) have energy use higher than today but have reached an equilibrium. The remaining growth scenarios feature several different rates, as low as 0.12% per year (S1) and up to 0.98% per year (S9). The idea of growth remains possible in our scenario set, but the rates of growth are lower than those considered by Kardashev (1964) for three of the scenarios (S1, S2, S6). The idea of slower-growth rates for future civilizational trajectories is consistent with insights from previous studies that have suggested rapid growth may be unsustainable, even across the solar system or beyond (e.g., Von Hoerner, 1975; Newman and Sagan, 1981; Haqq-Misra and Baum, 2009; Mullan and Haqq-Misra, 2019). For the future of human civilization, our scenario set provides several possibilities in which zero-growth (S3, S5, S10) or slow-growth (S1, S2, S6) is reached; these are alternative possibilities to the extremes of collapse (S4, S7, S8) or rapid growth (S9).
@@
Each of our scenarios represents a different future trajectory based on different assumptions about human social systems and technological capabilities. These assumptions include not only the physical infrastructure of civilization but also the values, worldviews, and myths/metaphors that form the basis of thinking and decisions in each scenario. The projection envisioned by Kardashev (1964) represents one possible future with a particular underlying myth/metaphor. Using the framing of Kardashev (1964), the appropriate description would be the myth of “inevitable growth,” which reflects other concerns that were prevalent on Earth during the 1960’s about limits to growth. Using the narrative in scenario S9, we describe this future as the myth of “Deus Ex Machina” to reflect the emergence of an advanced and self-directed technosphere. But other myths remain possible for the future (Table 5), and these other myths drive other futures, which in turn lead to different technosignatures. In thinking about the future of civilization on Earth, an important conclusion is that myths and worldviews can evolve with time, so the idea of continued growth into the long-term future is not necessarily an inevitability. For the search for technosignatures, an important conclusion is that the idea of rapid growth across the galaxy may be plausible but does not represent the only possible set of myths or worldviews that could sustain a long-lived technological civilization.

### v2

@@ A Relationship to Biosphere B Spatial Distribution C Development D Highest Order E Smallest Scale
As we did in §2.1, we conduct a cross-consistency assessment for the value pairs in the technology factors matrix to eliminate any pairs that would be inconsistent. The use of the LLM is even more helpful in expressing value pairs in natural language and providing explicit concise arguments for and against cross-consistency for this second matrix, as there are 113 unique possible combinations of value pairs in Table 3; however, the reasoning and assessment are tasks for the human researchers, despite the LLM’s natural language assistance, so we still need to manually inspect and evaluate all the initial results and rationales and make corrections. We again use a 4-point Likert scale to evaluate each value pair’s consistency; for our analysis of technology factors, we apply a more strict criterion and only maintain pairs of high consistency (with a score of 3) for propagation to the next steps of the analysis, while the remaining value pairs are rejected. The resulting technology cross-consistency matrix eliminates 50 of the 113 possible value pairs as inconsistent or unlikely, as shown in Figure 2.
Figure 2: Technology factors cross-consistency assessment. The 63 green cells with scores of 3 indicate consistent value pairs. The remaining 50 cells with scores less than 3 indicate inconsistent value pairs.
With the cross-consistency assessment complete, the remaining value pairs in the technology factors cross-consistency matrix (Fig. 2) result in eleven viable scenarios, which are summarized in Table 4. To better imagine and comprehend these value combinations, we again prompt the Claude LLM to express in natural language the value combinations of these eleven technology factors scenarios for a thousand years into the future, generating illustrative descriptions. The LLM scenario descriptions suggested that some scenarios show significant similarities to the extent that they could be grouped together. We choose to manually cluster the technology factors scenarios in an effort to further reduce the total number required for analysis. Afterwards, leveraging causal layered analysis as before, we inductively assign a myth/metaphor for each Cluster, again without assistance from the LLM.
@@ 2.3 Final Scenarios
We perform a compatibility assessment of the 33 viable global and technology scenarios to identify any combinations of global factors and technology factors parts that are inconsistent on the basis of their corresponding myths/metaphors that reflect their deeper civilizational foundations. We take an interactional theory stance (Friedman and Hendry, 2019) on the immaterial and material manifestations of the future anthroposphere, captured by the global and technology factors: we assume that they must be mutually shaped and compatible from a civilizational perspective at their deepest cultural level, that is, at the level of collective metaphors and myths. As before, we use the Claude LLM only to generate illustrative descriptions of the potential combinations of global factors and technology factors parts to help us in our imagining of the complexity of those anthroposphere combinations, but the assessment itself is conducted manually by the human researchers explicitly adopting an interactional theory stance. As discussed previously, the criterion used for this assessment is the compatibility of the deeper civilizational undercurrents of the global factors and technology factors parts of each candidate scenario as co-evolving and co-shaping aspects of the anthroposphere, especially as reflected in the level of their underlying myths/metaphors. This part of the assessment assigns only binary options for compatibility (yes/no) and does not include any additional scoring. The results of the compatibility assessment are shown in Figure 3, which leaves 16 remaining combinations between the global and technology factors that are compatible at the myth/metaphor level of their underpinning civilizations.
Figure 3: Compatibility assessment between global factors scenarios and technology factors scenarios. The 16 green cells marked ‘Y’ indicate consistent parings. The remaining 17 cells marked ‘N’ indicate inconsistent pairings.
@@ 5.1 Kardashev Scale Revisited
Technology exists to solve human biological and social problems, and technospheres evolve as the aggregate of these solutions. The methodology in this paper provides a way to link insights from the social sciences and humanities for understanding the complex dynamics of human societies into projections of the future that can inform the search for technosignatures. The underlying premise of this work is that human technology emerges from human needs and becomes intertwined with humans, which means that we cannot make meaningful projections about future technology without also thinking about how future developments would co-evolve with our human needs. This is one of the key insights missing from the projections by Kardashev (1964). We again note that we focus on the original arguments advanced by Kardashev (1964) because of the wide-ranging applications of these ideas in technosignature science, futures studies, and other applications. In many cases, the Kardashev scale is taken as a point of reference in the search for technosignatures (see, e.g., Sagan, 1973; Carrigan Jr, 2012; Cirkovic, 2015; Besteiro, 2019; Wright et al., 2022; Gray, 2020) as well as in future projections of human civilization (see, e.g., Baker, 2020; Namboodiripad and Nimal, 2021; Jiang et al., 2022; Zhang et al., 2023). Although many of these studies also offer critiques of the Kardashev scale, the common theme underlying all of these applications of Kardashev’s ideas is the assumption of continuous exponential growth. We therefore proceed with our discussion by referring to the original publication by Kardashev (1964), but the application of this discussion can be extended to contemporary studies that invoke continuous exponential growth for projecting possible technosignatures or trajectories of Earth’s future.
@@
We use the values in Table 9 to show approximate trajectories for energy use across the 1000-year timeline for all our scenarios in Figure 10. We draw a line or curve from the present-day energy use to the final value at the end of the timeline for each scenario. For some scenarios the value of energy use is lower than today (S4, S7, S8), which results from a prior period of growth before collapsing into the current state. Scenario S8 has a long-period oscillation with rapid growth followed by a collapse, currently midway through its third phase of growth. The three stable scenarios (S3, S5, S10) have energy use higher than today but have reached an equilibrium. The remaining growth scenarios feature several different rates, as low as 0.12% per year (S1) and up to 0.98% per year (S9). The idea of growth remains possible in our scenario set, but the rates of growth are lower than those considered by Kardashev (1964) for three of the scenarios (S1, S2, S6). The idea of slower-growth rates for future civilizational trajectories is consistent with insights from previous studies that have suggested rapid growth may be unsustainable, even across the solar system or beyond (e.g., Von Hoerner, 1975; Newman and Sagan, 1981; Haqq-Misra and Baum, 2009; Mullan and Haqq-Misra, 2019; Likavčan, 2024). For the future of human civilization, our scenario set provides several possibilities in which zero-growth (S3, S5, S10) or slow-growth (S1, S2, S6) is reached; these are alternative possibilities to the extremes of collapse (S4, S7, S8) or rapid growth (S9).
@@
Each of our scenarios represents a different future trajectory based on different assumptions about human social systems and technological capabilities. These assumptions include not only the physical infrastructure of civilization but also the values, worldviews, and myths/metaphors that form the basis of thinking and decisions in each scenario. The projection envisioned by Kardashev (1964) represents one possible future with a particular underlying myth/metaphor. Using the framing of Kardashev (1964), the appropriate description would be the myth of “inevitable growth,” which reflects other concerns that were prevalent on Earth during the early 1970s about limits to growth (e.g., Meadows et al., 1972). Using the narrative in scenario S9, we describe this future as the myth of “Deus Ex Machina” to reflect the emergence of an advanced and self-directed technosphere. But other myths remain possible for the future (Table 5), and these other myths drive other futures, which in turn lead to different technosignatures. In thinking about the future of civilization on Earth, an important conclusion is that myths and worldviews can evolve with time, so the idea of continued growth into the long-term future is not necessarily an inevitability. For the search for technosignatures, an important conclusion is that the idea of rapid growth across the galaxy may be plausible but does not represent the only possible set of myths or worldviews that could sustain a long-lived technological civilization.
@@ 5.3 Longevity of Technological Civilizations
5.4 Recommendations for Technosignature Detection
Further analysis of the scenario set developed in this study can be used to provide quantitative constraints on the search for technosignatures, such as by performing more sophisticated climate-chemistry modeling and then assessing the detectability of particular spectral features using instruments on current or future observatories. However, the results of the present analysis can still provide specific recommendations for advancing the search for technosignatures, in particular by re-examining concepts that are sometimes taken for granted in technosignature research. Three specific recommendations follow, based on the preceding discussion.
First is that increases in energy use may not be tightly coupled, as assumed in many applications of the Kardashev scale, and instead a technological civilization may prioritize exploration of its spatial scale over exploitation of its available energy (e.g., Berger-Tal et al., 2014). Even our rapid growth scenario S9 has only reached the energetic limit equivalent to a a Kardashev Type I civilization (i.e., using an energy equivalent to that available to Earth from incident solar radiation), but scenario S9 has a spatial scale that spans the entire solar system (which is the spatial scale associated with a Type II civilization). Technosignature searches that are predicated upon searching for Type I, II, or III civilizations in which the energetic and spatial scales are linked may find that no such civilizations exist at this precise limit, as all civilizations will require tradeoffs between prioritizing exploration and exploitation. Civilizations like Earth, and the scenarios considered in this study, may all be limited in their ability to harness luminous stellar energy and may not ever reach the theoretical limit of a Type I, II, or III civilization. Techosignature searches that seek to find evidence of a civilization that is approaching or even exceeding this limit would benefit from thinking more broadly about other possible metabolic mechanisms that prioritize exploitation over exploration. For example, the idea of a “stellivore” (Vidal, 2016) that harnesses energy by consuming stellar mass, rather than luminosity, is a theoretical possibility that would be observable as a compact accreting binary star system. Targeted searches of for such exotic technosignatures, such as analyses of compact accreting binaries (Vidal, 2019; Vidal et al., 2023; Vidal, 2024), provide an avenue for advancing technosignature searches beyond conventional approaches that look for Type I, II, or III civilizations.
Second is that spectral technosignatures may be degenerate or inconclusive when examining a single planet alone. The spectral characterization of terrestrial planets is becoming increasingly feasible and has motivated studies of spectral techosignatures that could conceivably be detectable with next-generation observatories (e.g., Kopparapu et al., 2021; Haqq-Misra et al., 2022b, a; Beatty, 2022); however, few, if any, of these spectral signatures would be conclusive evidence of extraterrestrial technology if observed in isolation. This is illustrated in our study with the degeneracy between the spectra of Earth in our technological scenarios (S5, S9, S10) and pre-agricultural Earth (R0). Breaking this degeneracy in our example would require searching for spectral technosignatures in the spectra of other planets (such as Mars or Venus), looking for artifact technosignatures in the outer parts of the system, or searching for communicative radio or laser signals. Targeted searches of planets of interest, such as planets thought to harbor liquid water or otherwise host habitable conditions, may be a necessary but insufficient condition for identifying the presence of a technosphere. In the best cases, the expansion of a technosphere to multiple planets in the system could provide corroborating evidence of extraterrestrial technology (such as a system that includes terraformed planets). In the worst cases, a planet of interest may harbor a technosphere that has no uniquely identifiable spectral signatures and no other detectable technosignatures. For such cases, exploration with remote spacecraft that can actually visit the planet may be the only way to detect evidence of such a technosphere. In general, efforts at spectral characterization of exoplanets should be conducted in tandem with other searches within the same system of interest in order to optimize the likelihood of finding evidence of a technosiganture.
Third is that technospheres similar to present-day or near-future Earth could conceivably be the most abundant technospheres in the galaxy or universe and thus the most likely to be detected. This could be because most technospheres are short-lived and do not ever reach the point of sustaining themselves across geologic timescales, but this could also be because long-lived technospheres are not so drastically different from Earth today. Nine of our scenarios (all but S1) could conceivably be sustained across geologic timescales without significant changes to their technospheres. Using Earth as a benchmark for technosignature detection studies may be a reasonable assumption to begin with; even though technology may be a relatively recent development in the history of Earth, this does not necessarily imply that future developments will necessarily lead to drastic shifts in the detectability of Earth’s technosphere. Our scenario set provides examples in which Earth’s spectral signature in the future is not so different from that of today, and it remains possible that the most abundant or detectable extraterrestrial technosignatures are comparable in magnitude to those on Earth today.